According to the Expectations Theory of interest rates, what would be the expected one-year rate, one year from today if the current two-year rate is 7.15% and the current one-year rate is 6.55%?

(Please provide all calculations, will leave a thumbs up if correct)
Answer #1

r(1,2) = 7.15%

r(1,1) = 6.55%

As per Expectation Theory,

(1 + r(1,2))2 = (1 + r(1,1))(1 + r(2,1))

(1.0715)2 = (1.0655)(1 + r(2,1))

r(2,1) = 7.75%

So,

One Year rate one year from today = 7.75%